OpenStudy (anonymous):

Solve the equation by completing the square -2x^2+x=0

OpenStudy (anonymous):

When completing the square 1) make sure x^2 have coeffient of 1, if not divide to get 1 2) Half and square the coefficient of the middle term 3)add the number you get to both side

OpenStudy (anonymous):

1) -2(x^2-1/2x)=0 2) -2(x^2-1/2x +1/16)=-2/16 -2(x^2-1/2x +1/16)=-1/8 -2 (x-1/4)^2=-1/8 -2(x-1/4)^2+ 1/8